Concentrated aqueous sulphuric acid is 98% H2SO4 by mass and has a density of 1.80 gmL . Volume of acid required to make 1 L of 0.1 M H2SO4 solution is

Grade:12th pass

1 Answers

Mohamed shakeel
5 years ago
98% of conc. h2so4 means100 g of h2so4 contains 98g of h2so4 by mass
molar mass of h2so4 is 98g/mol
no.of moles of h2so4 is 98g/98g/mol
                                = 1 mole
now volume of 100g of sulphuric acid =
volume=mass/density
         = 100/1.8
            = 55.55 mL
Now molarity = no.of moles of solute / volume of solution
M1= 1/55.55
now using the relation M1V1=M2V2
we have to find V1
M1=1/55.55
M2=0.1M
V2=1L
NOW
1/55.55 X V1 = 0.1 X 1
 WE WILL GET V1 as
V1 =5.55mL OR 5550 L